### v2

# G-DexGrasp: Generalizable Dexterous Grasping Synthesis Via Part-Aware Prior Retrieval and Prior-Assisted Generation

Abstract
Recent advances in dexterous grasping synthesis have demonstrated significant progress in producing reasonable and plausible grasps for many task purposes. But it remains challenging to generalize to unseen object categories and diverse task instructions. In this paper, we propose G-DexGrasp, a retrieval-augmented generation approach that can produce high-quality dexterous hand configurations for unseen object categories and language-based task instructions. The key is to retrieve generalizable grasping priors, including the fine-grained contact part and the affordance-related distribution of relevant grasping instances, for the following synthesis pipeline. Specifically, the fine-grained contact part and affordance act as generalizable guidance to infer reasonable grasping configurations for unseen objects with a generative model, while the relevant grasping distribution plays as regularization to guarantee the plausibility of synthesized grasps during the subsequent refinement optimization. Our comparison experiments validate the effectiveness of our key designs for generalization and demonstrate the remarkable performance against the existing approaches. Project page: https://g-dexgrasp.github.io/
Equal contribution Corresponding Authors: manyili@sdu.edu.cn, jianliu2006@gmail.com
1 Introduction
Figure 1: Given the diverse language-based task instructions and target objects of unseen categories, G-DexGrasp can synthesize high-quality dexterous grasps that can stably hold the object and are semantically aligned with the specified instruction.
Dexterous grasping synthesis  [13, 16, 34, 22, 40, 39, 41, 8, 48, 32] aims to generate reasonable and plausible solutions for the dexterous hands of robots in grasping and manipulating objects. A variety of works have proposed analysis-based  [40, 27, 6] or learning-based approaches  [11, 7, 24, 5, 41, 19] to synthesize dexterous hand grasps. However, they often focus on producing stable and diverse grasping solutions, but ignore the functionality goal of the grasps, i.e. whether the synthesized grasps can facilitate the following manipulation tasks with all kinds of objects.
To deal with various task intents and target objects, the researchers collect large amounts of annotations and learn to guide the synthesis of dexterous hand grasping. Some works [3, 35, 50, 14] classify the task intents into several categories and train conditional generative models to produce dexterous hand configurations. In parallel, some recent works  [4, 20, 42] utilize the powerful pre-trained large language models to generate natural language descriptions of the detailed grasping and synthesize hand configurations corresponding to the descriptions. These recent advances have shown significant progress in dealing with a variety of task purposes, but still struggle to effectively generalize to unseen object categories for task-driven grasping synthesis.
In this paper, we propose G-DexGrasp, a Generalizable retrieval-augmented Dexterous Grasping synthesis approach that can produce high-quality dexterous grasps for objects of unseen categories and unstructured task instructions, as shown in Figure 1. Considering the significant variations in object structures and shapes, the key is to retrieve generalizable grasping prior, which encompasses the fine-grained contact part as well as the affordance-related distribution of relevant grasping instances, aiming to assist the subsequent pipeline. The fine-grained contact part and affordance act as generalizable guidance to infer reasonable grasp configuration for unseen objects with a generative model, while the relevant grasping distribution plays as regularization to guarantee the plausibility of synthesized grasps during refinement optimization.
Our approach operates in three stages. The first stage analyzes the task instruction and target object to retrieve relevant grasping instances and organize the generalizable grasping prior. The second stage trains a generative model to estimate the coarse hand configuration, conditioned on fine-grained grasping guidance. The third stage takes the coarse estimation as initialization and optimizes to refine the parameters of dexterous hands regularized by the retrieved grasping prior. It is worth noting that our approach exploits the potential of training set to facilitate generalization without requiring extra labor-intensive joint-level annotations for more objects. Comparisons against existing approaches validates the effectiveness of our retrieval-augmented dexterous grasping synthesis. The ablation study confirms the effectiveness of our key designs in achieving remarkable generalization performance.
Our contributions are summarized as follows:
  • We present G-DexGrasp, a retrieval-augmented generation approach that can produce high-quality dexterous grasps for unseen object categories and task instructions.
  • We propose generalizable grasping prior, which acts as fine-grained guidance and semantic-aware regularization for the generalization of dexterous grasping synthesis.
  • We provide extensive experiments to validate the effectiveness of our key designs in achieving remarkable performance against existing approaches.
2 Related Work
2.1 Task-Agnostic Dexterous Grasping
Due to the high flexibility of dexterous hands, the primary challenge is to synthesize stable and plausible dexterous grasping solutions [32, 13, 16, 34, 22, 40, 39, 41, 8, 48]. Some works [13, 27, 40, 22] employ optimization algorithms and utilize grasping metrics as objective functions, such as force closure [6] for stability and penetration computation  [6, 40] for collision detection. For example, Obman [13] utilizes GraspIt!  [29] to produce feasible and collision-free grasps, while DexGraspNet [40] leverages an differentiable force closure estimator to generate stable grasp results. These approaches are able to produce a large variety of physically feasible grasps, but without considering the reasonable and natural arrangement of results, thus not convenient for the subsequent manipulation tasks in real scenarios.
On the other hand, data-driven approaches [11, 7, 24, 5, 41, 19, 47] have become popular, utilizing large annotated datasets to learn to generate plausible and human-like grasps. Many works [35, 16, 41, 20] train Variational Autoencoder networks (VAEs) to learn the grasp distribution and sample diversity grasp solution based on the given condition. Some others [28, 44, 4] utilize the recent diffusion model to sample dexterous grasps via the iterative denoising process. While data-driven approaches are capable of efficiently generating diverse and reasonable grasps, generating high-quality grasps remains a tough task.
To further enhance the grasping quality of the generated dexterous hands, some related works resort to intermediate representations to guide the grasping generation. Contact map [1] is a widely used representation with the vertex-wise contact information of the dexterous hands. Many approaches [49, 2, 10, 15, 18, 38, 26] generate reasonable hands based on the contact map prior information. In addition, FunctionalGrasp [51] uses touch codes [53], which record the contact status of each finger part w.r.t. the object’s surface. These fine-grained intermediate representations are used for a following conditional generation or optimization-based grasping refinement.
Figure 2: The three-stage pipeline of G-DexGrasp. (1) Generalizable grasping prior retrieval leverages pre-trained models to infer the fine-grained grasp configuration and retrieve relevant instances to form the grasping prior. (2) Part-aware grasping generation takes the target contact part and the retrieved prior as conditions to roughly initialize the hand. (3) Prior-guided grasping optimization takes the retrieved prior and other constraints as objective loss. It first optimizes the extrinsics only and then both extrinsic and intrinsic parameters.
2.2 Task-Oriented Dexterous Grasping
One of the main advantages of dexterous hands is their high flexibility for performing diverse manipulation tasks. Beyond ensuring stability and plausibility, increasing attention has been given to task-oriented dexterous grasping—i.e., generating grasps that facilitate subsequent manipulation.
Some related works [3, 35, 50] classify the task intents into several categories, such as “use", “pass", etc. They collect large amounts of grasping instances, including the task intent, target object, and dexterous hand pose, and train conditional generative networks to learn the distribution of task-oriented grasps. Moreover, AffordPose [14] collects the grasping instances and the fine-grained annotations, i.e. the part-level affordance labels such as “twist" and “handle-grasp". The affordance performs as a fine-grained classification of the task intent, thus producing more reasonable grasps for different objects and task intents.
The rapid development of natural language processing and the pre-trained large language models have enabled understanding and reasoning about task intent from natural language. Task2Grasp [4] uses the CLIP text encoder [31] to process the task instruction and trains a diffusion model conditioned on the text encoding and the object point cloud to generate hand poses. By contrast, GraspAsYouSay [42] utilizes more detailed, finger-level instructions to guide grasp generation. SemGrasp [20] fine-tunes the pre-trained Multimodal Large Language Model (MLLM), i.e. LLaVA [23], to predict grasp tokens based on the task instruction and object point cloud, which are then decoded into hand parameters via a pre-trained hierarchical VAE. The concurrent work AffordDexGrasp [43] generates generalizable affordances to assist grasping.
In light of the generalization of task-oriented dexterous grasping, the key is to infer the grasping arrangement for unseen objects w.r.t. the specified task instruction. Considering the large variation of object shapes and delicate grasping parameters, the performance of the above approaches is limited when directly transferred to deal with unseen objects. Inspired by the concept of generalizable parts [36, 9, 37] and the high-quality dexterous hand synthesis with part-level affordances [14], we utilize the pre-trained large models to retrieve the fine-grained generalizable grasping prior and synthesize the reliable and semantically aligned dexterous hands for unseen objects.
3 Method
3.1 Overview
Our problem follows the task-oriented dexterous grasping synthesis, where the input contains a task instruction $L$ in the form of natural language and a target object represented as 3D point cloud, i.e. $O\in\mathbb{R}^{N\times 3}$. The output is the pose parameters of the synthesized dexterous hand, denoted as $G=(T,R,\Theta)$. $T\in\mathbb{R}^{3}$ and $R\in SO(3)$ are the extrinsic parameters representing the transformation and rotation of the global hand, while $\Theta\in\mathbb{R}^{d}$ is the intrinsic parameters representing angle values for the joints of the dexterous hand. We have $d=16$ with the articulated kinematic MANO hand model [13, 33]. The synthesized hand should be able to stably and anthropomorphically hold the target object and be semantically aligned with the specified task instruction. In this paper, we especially focus on generalizable dexterous grasping synthesis. That is, the input task instructions and target objects are not limited within the distribution of the training data, but rather generalized to the diverse task instructions and unseen object categories.
We leverage the retrieval-augmented generation (RAG) strategy in our approach. To support the retrieval of relevant grasping instances, we pre-process the existing dataset as $\Omega=\{(L_{i},O_{i},G_{i},A_{i},P_{i},M_{i})\}_{1}^{N}$. Each data $\Omega_{i}$ includes the task instruction $L_{i}$, object model $O_{i}$, hand pose $G_{i}$, and the annotated affordance type $A_{i}$, contact part $P_{i}$ of the grasping instances, part segmentation $M_{i}$ of 3D objects.
Our approach consists of three stages. As shown in Figure 2, we first analyze the task instruction and target object to retrieve relevant grasping instances and organize a generalizable grasping prior, which includes representative hand intrinsics, contact map distribution, and the target’s contact part. Next, we train a part-aware generative network to estimate extrinsic hand parameters conditioned on the given intrinsics and contact part. Finally, using the retrieved intrinsics and generated extrinsics as the initial pose, we refine the hand via iterative optimization with physical constraints and the retrieved prior.
3.2 Generalizable Grasping Prior Retrieval
In this stage, given the task instruction and the target object, we utilize the pre-trained Multimodal Large Language Model (MLLM) to infer the task-oriented and fine-grained grasping arrangement, i.e., the affordance type and contact part. Then we retrieve relevant grasping instances from the dataset and organize their distributions as the grasping prior. The prior is generalizable because both the affordance and the contact part are generalizable across object categories. The former indicates the rough hand intrinsics and the latter refers to the grasping region on the object surface.
Task Instruction Analysis. We first prompt the MLLM to determine the fine-grained grasping arrangement. It takes the task instruction $L$ and the rendered image $I$ of object $O$, and outputs the affordance type $A$ and the name of contact part $P$. The image is rendered from the selected viewpoint to show its shape and appearance as much as possible. In our prompt, we provide a set of pre-defined affordance types as well as their descriptions and restrict the MLLM to select one from them. For example, given the image of a kettle and the instruction “could you please hold the kettle to pour some tea?", the MLLM should infer the affordance type “handle-grasp" and the contact part “handle".
Contact Part Localization. Given the inferred part name, we need to localize the corresponding part region of the 3D object. We employ the pre-trained GLIP model [21], which takes the rendered image and the contact part name as input and outputs the detected 2D bounding boxes of the contact part. Then the pixels within the bounding boxes are projected to the 3D point cloud of the object, with the help of an additional depth map rendered from the same viewpoint. Due to the occlusion with the rendered image, we obtain some partial 3D parts from the detected bounding boxes. In our implementation, we utilize the segmented objects in the pre-processed dataset and vote from these partial parts to find the complete contact part $P$. For objects without pre-segmentation, an alternative solution is to fuse multi-view segmentation using existing methods [25, 52].
Grasping Prior Retrieval. To retrieve the relevant grasping instances from the dataset, we define the dissimilarity between grasping instances $\Omega_{i}$ and $\Omega_{j}$:
$Dist(\Omega_{i},\Omega_{j})=D_{A}(A_{i},A_{j})+D_{P}(P_{i},P_{j})+D_{\Theta}(\Theta_{i},\Theta_{j}),$ (1)
where $D_{A}$, $D_{P}$, and $D_{\Theta}$ measure the dissimilarity between the affordance types, contact parts, and the hand intrinsics, respectively. With $OBB$ representing the sorted size of the contact part’s oriented bounding box, we have
$D_{A}(A_{i},$ $A_{j})=\begin{cases}\infty,A_{i}\neq A_{j}\\ 0,A_{i}=A_{j}\end{cases},\quad D_{\Theta}=\left\|\Theta_{i}-\Theta_{j}\right\|,$ (2)
$D_{P}(P_{i},P_{j})=||OBB(P_{i})-OBB(P_{j})||.$
Based on this similarity, we perform K-means clustering on the dataset. Each clustering contains the grasping instances that belong to the same affordance type and have similar contact part shapes and intrinsic hand poses.
Instead of the other retrieval-augmented generation approaches that only utilize the most similar instance, we organize the grasping knowledge of each cluster as a generalizable grasping prior $\Gamma=(P,\hat{\Theta},C)$ that includes the localized contact part $P$, the representative intrinsic pose $\hat{\Theta}$, and the contact map distribution $C$ for hand. The representative pose $\hat{\Theta}$ is the nearest neighbor instance to the mean of the clustering. The hand contact map distribution $C$ refers to the mean and standard deviation of the Gaussian distribution of the contact maps in the cluster, where a hand contact map reflects the nearest distance of each mesh vertex of the hand to the object surface [10]. Note that the computation of generalizable grasping prior is scalable, meaning that we can obtain convincing priors from a small-scale dataset and can flexibly extend it with more types of grasp instances.
3.3 Part-Aware Grasping Generation Network
The second stage aims to estimate the rough hand parameters $\hat{G}=(\hat{T},\hat{R},\hat{\Theta})$ based on the given object and the inferred grasping arrangement. We directly take the retrieved representative hand pose $\hat{\Theta}$ as the intrinsic parameters and train a network to predict corresponding extrinsic parameters, i.e. $\hat{T}$ and $\hat{R}$, in the coordinate system origined at contact part center. In order to generalize to unseen objects, we design the part-aware grasping generation network, supervised by the ground-truth hand extrinsics from the training set without intrinsic parameters.
The network is implemented as a conditional variational auto-encoder, as shown in Figure 3. The input condition includes contact part point cloud $O^{P}$, part size $S^{P}$, representative intrinsic parameters $\hat{\Theta}$, centroid $V$ of the rest of the object. The point cloud $O^{P}$ is normalized into a unit sphere with a scaling factor, while the part size $S^{P}$ is the axis-aligned bounding box size. The centroid $V$ is defined as the center location of the object excluding the contact part, w.r.t. the center of the contact part. It prevents a large overlap between the predicted hand and the object.
Figure 3: Network architecture of part-aware generation network.
To encode the input conditions, for the part point cloud, we concatenate each point with its normal vector and the scaling factor, and use PointNet++ [30] to produce a $128$-dimensional latent vector. The representative intrinsic parameters $\hat{\Theta}$ are represented as the one-hot vector of its corresponding cluster in the dataset. This one-hot vector, centroid $V$ and part size $S^{P}$ are encoded as $128$-dimensional latent vectors using linear layers, respectively. The latent vectors are concatenated to compose the latent condition vector. The CVAE network takes the condition vector and learns to reconstruct extrinsic parameters, i.e. $T$ and $R$, which are fed into a differentiable MANO layer [13] to produce the mesh of the corresponding dexterous hand.
The loss function of the grasping generation network is
$\mathcal{L}=\mathcal{\lambda}_{KLD}\mathcal{L}_{KLD}+\mathcal{L}_{recon}+\mathcal{\lambda}_{pen}\mathcal{L}_{pen},$ (3)
where $\mathcal{L}_{KLD}$ is the KL-divergence loss of the CVAE network to align the learned distribution with a Gaussian distribution. The reconstruction loss $\mathcal{L}_{recon}$ is the mean square loss between the reconstructed hand and the ground-truth
$\mathcal{L}_{recon}=\mathcal{\lambda}_{V}\textperiodcentered||{\hat{\mathcal{V}}}-\mathcal{V}||+\mathcal{\lambda}_{R}\textperiodcentered||\hat{R}-R||+\mathcal{\lambda}_{T}\textperiodcentered||\hat{T}-T||,$ (4)
where $\hat{\mathcal{V}}$, $\hat{R}$, $\hat{T}$ represent the predicted mesh vertices, global rotation, and global translation of the hand, while $\mathcal{V}$, $R$, $T$ are the ground-truth. The penetration loss $\mathcal{L}_{pen}$ penalizes the overlapping between the dexterous hand and the object
$\mathcal{L}_{pen}=\sum_{i}\max(0,(O_{i}-\hat{\mathcal{V}}_{i})\cdot\mathbf{n}^{O}_{i}-c_{pen}),$ (5)
where $O_{i}$ and $\mathbf{n}^{O}_{i}$ denote the object mesh vertex and its corresponding normal vector, $\hat{\mathcal{V}}_{i}$ is the nearest hand vertex to $O_{i}$, $c_{pen}$ is a pre-determined threshold.
3.4 Prior-Guided Grasping Optimization
It is difficult to have a network to directly generate precise grasping parameters, especially for unseen object categories. Therefore, we use the retrieved generalizable grasping prior as the guidance of a differentiable grasp optimizer to refine the roughly estimated hand parameters.
Given the initial dexterous hand parameters $\hat{G}=(\hat{T},\hat{R},\hat{\Theta})$, we use the differentiable MANO layer to obtain the hand mesh and compute the objective function. The objective function consists of a variety of terms from three aspects. One is the retrieved grasping prior guidance, which is composed of $\mathcal{L}_{contact}$ to encourage the hand to contact with the target part and $\mathcal{L}_{cmap}$ to have the contact map conforming to the retrieved distribution. We have
$\mathcal{L}_{contact}=\frac{1}{n}\sum_{j=1}^{n}C_{j}\underset{i\in N}{\min}\|\ \hat{\mathcal{V}}_{j}-O^{P}_{i}\|,$ (6)
$\mathcal{L}_{cmap}=\frac{1}{n}\sum_{j=1}^{n}e^{-\sigma_{j}}\|\hat{C}-C\|,$
where $C$ is the averaged hand contact map of the retrieved grasp cluster and $C_{j}$ is the averaged contact map value of the $j$th hand vertex. $\sigma_{j}$ is the standard deviation of the contact maps of the retrieved grasp cluster at the $j$-th vertex. $\hat{C}$ is the contact map of the predicted hand. $\hat{\mathcal{V}}_{j}$ and $O^{P}_{i}$ are the vertices of the hand and contact part. Following the prior works, the second is the physical feasibility constraints of the grasping, including the differentiable force closure term $\mathcal{L}_{fc}$ [40] to enhance the grasping stability, the penetration term $\mathcal{L}_{pen}$ [10] (Eq. 7) and $\mathcal{L}_{revpen}$ to avoid hand-object overlapping, i.e.
$\mathcal{L}_{revpen}=\sum_{i}\max(0,(\hat{\mathcal{V}_{i}}-O_{i})\cdot\mathbf{n}^{\hat{\mathcal{V}}}_{i}-c_{revpen}),$ (7)
where $\hat{\mathcal{V}_{i}}$ and $\mathbf{n}^{\hat{\mathcal{V}}}_{i}$ denote the predicted hand mesh vertex and its normal vector, $O_{i}$ is the nearest object vertex to $\hat{\mathcal{V}_{i}}$, $c_{revpen}$ is a pre-determined threshold. The third is the human-like grasping constraints, containing $\mathcal{L}_{joint}$ [40] to restrict each angle within the range of human joint and $\mathcal{L}_{spen}$ [53] to avoid self-penetration of the hand.
Our optimization contains two sub-processes. Specifically, the first aims to refine the extrinsic parameters while preserving the intrinsic hand pose. It only utilizes the retrieved grasping prior guidance as the objective function. The goal is to quickly move the hand to its target position and orientation, rather than struggling at an incorrect local minimum. The second stage fine-tunes all the parameters to obtain the final dexterous hand $G=(T,R,\Theta)$, where all the above terms are combined with their weights as the objective function. We use Adam optimizer for the two processes, where the first uses a learning rate of 0.005 and the second 0.001. Each stage contains 700 iterations of updates.
4 Experiments
Experiments Quality and Stability Semantic Alignment
[…]
AffordPose [14] 6.86/10.55 0.96/1.21 10.66/9.93 81.0/42.5 3.84/1.72
GrabNet [35] 8.76/11.07 1.08/1.47 10.58/9.62 75.6/45.5 3.67/1.63
GraspTTA [15] 5.34/7.98 0.89/0.92 8.17/7.61 74.3/46.5 2.03/0.88
Ours 3.13/2.94 0.82/0.29 5.39/4.27 82.6/71.6 3.72/3.93
Ours 3.18/2.04 0.87/0.22 4.93/4.34 89.74/96.3 3.94/4.44
Table 1: Quantitative comparison on the seen///unseen object categories. GrabNet and GraspTTA use BERT to encode task instructions. Ours uses pre-trained models to infer and localize contact parts. Ours uses ground-truth part segmentation with inferred contact part selection. Best and second-best results are marked in red and green, averaged over seen and unseen cases.
4.1 Experiment Settings
Dataset. We use the AffordPose dataset [14], which provides dexterous hand data, 3D object shapes, and part-level affordance labels for each grasp. It includes 8 interaction types across 13 object categories. We select five grasp-related interactions—handle-grip, twist, wrap-around-grip, base-support, and trigger-squeeze—and split the data 8:1:1 for training, validation, and testing. The training set is used for retrieval and generation network training, while the test set validates performance on seen categories.
We select novel objects to test generalization, including 92 objects from 11 categories in OakInk [50] (e.g., frying pan, bowl), 81 objects from 13 categories in PartNet-Mobility [46] (e.g., suitcase, coffee machine), and 126 objects from 14 categories in 3D-Net [45] (e.g., broom, umbrella), forming our unseen test set.
[…]
Implementation. We invoke GPT-4o as the pre-trained MLLM to infer the affordance type and contact part of the dexterous grasping. We sample $N=2048$ points from the object surface as the input object points. During training, we use the Adam optimizer with a learning rate of 1e-3 to train the network for 400 epochs. The loss weights are $\mathcal{\lambda}_{V}=100$, $\mathcal{\lambda}_{R}=5$, $\mathcal{\lambda}_{T}=30$, $\mathcal{\lambda}_{KLD}=10$, $\mathcal{\lambda}_{pen}=1$.
Figure 4: Qualitative comparison. Given object from unseen categories and language-based task instruction, our approach exhibits remarkable performance in the generalization. Note that the contact parts are highlighted in the first row for visualization, not as input.
4.2 Comparisons
We compare against the state-of-the-art methods, which focus on taking the affordance type labels or the embedding of textual task descriptions as conditions for dexterous grasping synthesis. Note that AffordPose [14] and OakInk [50] actually use the same GrabNet network [35] but training data with different sets of task labels. So we select AffordPose as their representative. In addition, with the absence of released code of Text2Grasp [4], we follow their method which uses pre-trained BERT [17] to encode the task descriptions and select the existing network architectures GrabNet [35] and GraspTTA [15] to examine the performance on task-oriented dexterous grasping, denoted as GrabNet and GrabTTA. We also provide two different versions of our approaches. One is our full method which uses the pre-trained model to localize contact parts, while the other, denoted as Ours*, uses the ground-truth part segmentation with inferred contact part selection.
The qualitative and quantitative evaluations are reported in Table 1 and Figure 4. The first two rows of Table 1 use the same network architecture but different conditions, i.e. object point cloud and either the affordance type (first row) or the task description (second row). It shows that the task descriptions act better than affordance types in guiding the dexterous hand to contact the correct region on the unseen objects in our test set (see the semantic alignment metrics). Due to the same network architecture, they exhibit similar performance in terms of grasping quality and stability. By contrast, GraspTTA shows better stability but similar semantic alignment with GraspNet, due to the same task description condition but a more advanced network.
However, more importantly, our approach achieves a remarkable improvement in the performance on unseen object categories, thanks to our retrieval-augmented grasping generation strategy. In addition, when utilizing the ground-truth part segmentations, i.e. ours*, all the metrics exhibit a further improvement, especially the semantic alignment. It adequately validates the transferability of our generalizable grasping prior on unseen object categories and the potential of our approach with the further development of open-vocabulary part segmentation approaches.
4.3 Ablation Study
We validate our key designs, which are of vital importance to correctly and efficiently exploit the retrieved information for dexterous hand generation. Qualitative and quantitative evaluations are reported in Figure 5 and Table 2.
Experiments Quality and Stability Semantic Alignment
[…]
Object-based Net. 10.06 0.66 8.64 51.17 1.43
Part Rand. Init. 1.34 0.21 6.40 57.85 1.04
W/O Optim. 17.43 1.55 3.3 66.88 1.92
W/O Prior Guid. 9.13 0.82 3.65 68.22 2.30
One-Stage Optim. 3.18 0.28 5.02 69.23 2.51
Ours 2.94 0.29 4.27 71.6 3.93
Table 2: Quantitative evaluation of our ablation study to validate the importance of the retrieved prior in the grasping synthesis pipeline.
[…]
Figure 5: Qualitative evaluation of ablation study. Compared to our results, (b) and (c) validate the importance of using a part-aware network for hand initialization, while (d), (e), and (f) reveal the effectiveness of the prior-guided optimization to produce stable grasps.
How does the retrieved information assist the grasping generation? We design two different variants of our part-aware grasping generation network, named as Object-based Net. and Part Rand. Init. The former uses an object-conditioned network and the latter uses a random initialization around the target part, to replace our part-aware network which generates initial extrinsic parameters for the subsequent optimization. Specifically, compared to our full approach, the object-conditioned network removes the part-related conditions, i.e. $V^{*}$ and $S^{P^{*}}$, and uses the object point cloud to replace part point cloud $O^{P^{*}}$. On the other hand, for Part Rand. Init., we follow the method of DexGraspNet [40] to initialize the extrinsic parameters to place the dexterous hands around the contact part.
The first two rows of Table 2 report the results. We can see that the object-based network performs much worse on all the metrics. Because the terrible extrinsic initializations often cause dexterous hands with heavy penetration with the object, making it hard for the following optimization to refine them to stable and reasonable grasps. By contrast, part-aware random initialization heuristically places the dexterous hands around the contact part, thus resulting in good performance in terms of quality and stability. However, the randomness affects the robustness of the approach, causing slightly worse simulation and part accuracy. In summary, these baselines validate the necessity of using a part-aware network to initialize extrinsic parameters.
How does the retrieved information improve the refinement grasping optimization? We conduct the three alternative options w.r.t. our prior-guided grasping optimization. The experiment W/O Optim. removes the optimization and directly takes the initialized dexterous hands as the final outputs. The experiment W/O Prior Guid. uses the same retrieval-assisted dexterous hand initialization as our full approach, but doesn’t use any retrieved information during the optimization process. More specifically, it removes $\mathcal{L}_{cmap}$ and replace the $\mathcal{L}_{contact}$ with the $\mathcal{L}_{close}$ term in [12], which only encourages the fingers of dexterous hand to contact with the object without any information from the retrieved contact maps. The experiment One-Stage Optim. exploits the retrieved prior as we did, but optimizes both the intrinsic and extrinsic parameters in one single stage.
The third to the fifth row of Table 2 show their results. Without optimization, although the hands often approach the correct parts, they exhibit heavy penetration with the object and cannot hold the object stably. Without the prior guidance in optimization, it often produces unnatural grasps with larger penetration and sometimes slide to contact with other parts. By contrast, the one-stage optimization can produce good results for many cases, but would fail when the initialized extrinsic parameters cause opposite rotations or large distances to the object. In summary, these baselines validate the necessity of leveraging the retrieved prior as the objective of the optimization to guarantee the quality and semantic alignment of the refined dexterous grasps.
5 Conclusion
We present G-DexGrasp which exploits the retrieval-augmented generation for dexterous grasping synthesis on unseen object categories and diverse task instructions. The key is to retrieve the generalizable grasping prior and take it as the fine-grained guidance for the subsequent generation and refinement optimization. Through extensive experiments, we validate the importance of leveraging a part-aware network to generate reasonable rough dexterous hands, and taking the retrieved prior as objective in a two-stage refinement optimization. Our results exhibit remarkable performance in terms of grasping quality and stability, as well as semantic alignment with task instructions.
However, our approach requires an open-vocabulary part segmentation for the unseen objects, which, though not the focus of this work, limits the performance of the synthesized dexterous hands. In addition, extending the pre-processed dataset with a wider range of task-oriented hand-object interaction instances, especially for those tasks demanding delicate dexterous hand postures, would further stimulate more manipulation tasks in various scenarios.
Acknowledgments.
This work was supported by National Natural Science Foundation of China (62272082, 62302269, 62027826, 62322207), the Excellent Young Scientists Fund Program (Overseas) of Shandong Province (2023HWYQ-034), Shenzhen University Natural Sciences 2035 Program (2022C007), and the joint Fund General Project of Liaoning Provincial Department of Science and Technology (2024-MSLH-352).